# Data Imputation by Pursuing Better Classification: A Supervised Kernel-Based Method

### v2

@@ 5 Conclusion
This paper proposed a novel two-stage data imputation framework, aiming to optimize the similarity relationships between data in order to guide the completion of missing features by pursuing better classification accuracy. In the first stage, we unify the tasks of kernel matrix imputation and classification within a single framework, enabling mutual guidance between the two tasks in an alternating optimization process to improve similarity relationships. The introduction of a perturbation variable enhances the robustness of the algorithm’s predictions. In the second stage, we achieve, for the first time, the recovery of data features from a given kernel matrix, effectively utilizing the optimized information obtained in the first stage. By leveraging the supervision information through two stages, we have obtained a more flexible approach for data imputation, which provides significant advantages when dealing with high missing data rates. Numerical experiments have validated that our algorithm achieves higher prediction accuracy and demonstrates more stable performance compared to other methods on the test data. In future research, extending the ideas of this study to neural networks would be a fascinating endeavor. This expansion would provide us with a way to handle even larger-scale data. It is important to note that neural networks typically operate in over-parameterized regimes, leveraging their high model complexity to estimate labels effectively. However, this also makes it challenging to control the impact of supervised information on the results of data imputation in the absence of appropriate regularization terms.
@@ Appendix A Decoupling of Various Kernel Function Computations
    Linear Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\boldsymbol{x}^{\top}\boldsymbol{y}$,
@@
    Polynomial Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\left(\boldsymbol{x}^{\top}\boldsymbol{y}+r\right)^{d}\ (r\geq 0\ \mathrm{and}\ d\geq 1)$,
@@
    Gaussian Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\mathrm{exp}\left(-\gamma\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2}\right)$,
@@
    Laplacian Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\mathrm{exp}\left(-\gamma\|\boldsymbol{x}-\boldsymbol{y}\|_{1}\right)$,
@@
    Sigmoid Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\mathrm{tanh}\left(\gamma\boldsymbol{x}^{\top}\boldsymbol{y}+r\right)\ (\gamma>0\ \mathrm{and}\ r\geq 0)$,
@@
    $\chi^{2}$ Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\sum_{p=1}^{d}\frac{2x_{p}y_{p}}{x_{p}+y_{p}}$,

### v3

@@ 5 Conclusion
This paper proposed a novel two-stage data imputation framework, aiming to optimize the similarity relationships between data to guide the completion of missing features by pursuing better classification accuracy. In the first stage, we unify the tasks of kernel matrix imputation and classification within a single framework, enabling mutual guidance between the two tasks in an alternating optimization process to improve similarity relationships. The introduction of a perturbation variable enhances the robustness of the algorithm’s predictions. In the second stage, we achieve, for the first time, the recovery of data features from a given kernel matrix, effectively utilizing the optimized information obtained in the first stage. By leveraging the supervision information through two stages, we have obtained a more flexible approach for data imputation, which provides significant advantages when dealing with high missing rates. Numerical experiments have validated that our algorithm achieves higher prediction accuracy and demonstrates more stable performance compared to other methods. In future research, extending the ideas of this study to neural networks would be a fascinating endeavor. This expansion would provide us with a way to handle even larger-scale data. It is important to note that neural networks typically operate in over-parameterized regimes, leveraging their high model complexity to estimate labels effectively. However, this also makes it challenging to control the impact of supervised information on the results of data imputation in the absence of appropriate regularization terms.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (62376155). The authors would like to thank the anonymous AE and reviewers for their insightful comments.
@@ Appendix A Decoupling of Various Kernel Function Computations
    Linear Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\boldsymbol{x}^{\top}\boldsymbol{y}$, and
@@
    Polynomial Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\left(\boldsymbol{x}^{\top}\boldsymbol{y}+r\right)^{d}\ (r\geq 0\ \mathrm{and}\ d\geq 1)$, and
@@
    Gaussian Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\mathrm{exp}\left(-\gamma\|\boldsymbol{x}-\boldsymbol{y}\|_{2}^{2}\right)$, and
@@
    Laplacian Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\mathrm{exp}\left(-\gamma\|\boldsymbol{x}-\boldsymbol{y}\|_{1}\right)$, and
@@
    Sigmoid Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\mathrm{tanh}\left(\gamma\boldsymbol{x}^{\top}\boldsymbol{y}+r\right)\ (\gamma>0\ \mathrm{and}\ r\geq 0)$, and
@@
    $\chi^{2}$ Kernel: $k(\boldsymbol{x},\boldsymbol{y})=\sum_{p=1}^{d}\frac{2x_{p}y_{p}}{x_{p}+y_{p}}$, and